Four rowers weighing 152, 156, 160, and 164 pounds make up a rowing team. Calculate the sample mean for each of the possible random samples with a size two replacement. They can be used to calculate the sample mean's probability distribution, mean, and standard deviation.



1
Expert's answer
2022-05-03T11:01:32-0400

We have population values 152, 156, 160, and 164 population size N=4 and sample size n=2.

Mean of population "(\mu)" = 

"\dfrac{152+156+160+164}{4}=158"


Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{N}=20"


"\sigma=\sqrt{\sigma^2}=\sqrt{20}=2\sqrt{5}"

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{4}C_2=6."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 152, 156 & 154 \\
 \hdashline
 2 & 152, 160 & 156 \\
 \hdashline
 3 & 152, 164 & 158 \\
 \hdashline
 4 & 156, 160 & 158 \\
 \hdashline
 5 & 156, 164 & 160 \\
 \hdashline
 6 & 160,164 & 162 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) &\bar{X}^2 f(\bar{X})\\ \hline
 154 & 1/6 & 154/6 & 23716/6\\
 \hdashline
 156 & 1/6 & 156/6 & 24336/6\\
 \hdashline
 158 & 2/6 & 316/6 & 49928/6\\
 \hdashline
160 & 1/6 & 160/6 &25600/6\\
 \hdashline
162 & 1/6 & 162/6 & 26244/6\\
 \hdashline
\end{array}"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=158=\mu"


The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{149824}{6}-(158)^2=\dfrac{40}{6}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{\dfrac{20}{3}}\approx2.5820"